OpenStudy (anonymous):

Im having trouble understanding what this question is asking for exactly? Can anyone help explain? If a function, f(x) is shifted to the left four units, what will the transformed function look like?

OpenStudy (solomonzelman):

what is your f(x), (your function), is it given to you?

OpenStudy (anonymous):

It doesn't give you a function that's the thing, it just says "if a function" in general.

OpenStudy (solomonzelman):

Ohh, so I will give you a couple of examples, of a shift in a function. \(\large\color{ black }{\large {\bbox[5pt, lightyellow ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ \color{blue}{ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x \normalsize\color{red}{ +~\rm{c}} } } &~\rm{\color{blue}{ to~~the~~left ~}} \\ \text{ } \\ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x} \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x} \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\) \(\large\color{ black }{\large {\bbox[5pt, lightcyan ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ -~\rm{c} })^2 &~\rm{to~~the~~right~} \\ \text{ } \\ \color{blue}{ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ +~\rm{c}} }\color{blue}{)^2}&~\rm{ \color{blue}{to~~the~~left} ~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

you are adding "c" for a shift of c units to the left, so to speak, "inside the parenthesis" see?

OpenStudy (solomonzelman):

So they are asking you to say, (I think) that you would add +4 into every x. YOu know what I mean, right. It will be f(x+4), yes?

OpenStudy (anonymous):

That seems to make sense from the video I was watching. I didn't know if that's what they were asking for or not though. So I would just right f(x)+4?

OpenStudy (solomonzelman):

yes. I think...